Help!
If P(A or B) = 0.9, P(A) = 0.7 and P(B) = 0.5, determine P(A and B).
Thanks!
Melissa

Answer by sudhanshu_kmr(1152)


P(A U B) = P(A) + P(B) - P(A n B)
using this formula...
=> 0.9 = 0.7 + 0.5 - P(A n B)
=> P(A n B) = 1.2 - 0.9 = 0.3

SO, P( A AND B) = 0.3